OpenStudy (anonymous):

derivation for attenuation constant and phase constant for propagation constant of plane wave in lossy media

OpenStudy (anonymous):

i'd done some equating stuffs and got \[\alpha ^{2}-\beta^{2}+j2\alpha \beta=-w^{2}\mu \epsilon'+jw^{2}\mu \epsilon''\] now i just need to equate the real and imaginary parts which will be\[\alpha^{2}-\beta^{2}=-w^2 \mu \epsilon'\]and\[2 \alpha \beta=w^2 \mu \epsilon''\]

OpenStudy (anonymous):

but i cant reach the final derived expression,\[\alpha=w \sqrt{\frac{\mu \epsilon'}{2}[\sqrt{1+(\frac{\epsilon''}{\epsilon'})^{2}}-1]}\]\[\beta=w \sqrt{\frac{\mu \epsilon'}{2}[\sqrt{1+(\frac{\epsilon''}{\epsilon'})^{2}}+1]}\]

OpenStudy (anonymous):

its like after equating the real and imaginary parts, the alpha and beta cant be factorized ompltely

OpenStudy (nincompoop):

use wolframalpha and check

OpenStudy (anonymous):

wolfrom allows so many unknowns?

OpenStudy (nincompoop):

it can

OpenStudy (anonymous):

how to make it interpret correctly? i tried solve a^2-b^2+j(2ab)= -(omega)^2 (mu)(epsilon)+j(omega)^2(mu)(sigma/omega) but it doesnt return the answer i need

hartnn (hartnn):

isolate b from 2ab = ... put it in a^2-b^2 = .... you will get something in b^2 and b^4 assume b^2 =x so you will get quadratic in x use quadratic formula to solve then plug back x = b^2 and take square root to find b (cumbersome process, but you should get that)